OpenStudy (anonymous):

Which of the following correctly represents the coordinates of the foci of the ellipse shown below? (x-7)^2/4+(y+3)^2/16=1 A. (7±2√3, -3) B. (7, -3±2√5) C. (7±2√5, -3) D. (7, -3±2√3) am i correct ?? its D i just want to really be sure ...

OpenStudy (tkhunny):

Center: Eyeball (7,-3) Half-Length of Major Axis: Eyeball a = 4 (y-direction) Major Foci: (7,1), (7,-7) -- This eliminates A and C, above. The foci must be x = 7. Half-Length of Minor Axis: Eyeball b = 2 (x-direction) Minor Foci: (9,-3), (5,-3) Solve for c^2 = a^2 - b^2 = 16 - 4 = 12, so c = \(2\sqrt{3}\) -- This eliminates B.

OpenStudy (tkhunny):

You just have to wade through it - one piece at a time. Some things are obvious. Others are not.

OpenStudy (anonymous):

thank you n_n @tkhunny

OpenStudy (anonymous):

this really does reassure me n_n

OpenStudy (tkhunny):

Next time, you show your work and don't make me do it all.

OpenStudy (anonymous):

ohhh okk ill change strategies --- ill state my answer and workn_n to go along with my reasoning

OpenStudy (tkhunny):

Good. If you just say what answer you managed, how does anyone know you didn't just copy it out of the back of the book? Show your work. It makes a difference and makes everything better. :-)